OpenStudy (anonymous):

√c (3 + 2√5) and √6 (7√3 + 6) Please help???

OpenStudy (anonymous):

is that a c in front?

OpenStudy (anonymous):

its asking for the square root of 3 then(3+2√5) and then the other question

OpenStudy (anonymous):

@ satellite

OpenStudy (anonymous):

The left :/

OpenStudy (anonymous):

):

OpenStudy (anonymous):

What are we suppose to do here?

OpenStudy (anonymous):

Multiply and Simplify

OpenStudy (anonymous):

Multiply/Distribute in this case, and simplify

OpenStudy (anonymous):

I got the first one, just need help with the second

OpenStudy (anonymous):

Y dont you retype the question?

OpenStudy (anonymous):

It isnt so clear

myininaya (myininaya):

\[\sqrt{3} (3+2 \sqrt{5}) \sqrt{6}(7 \sqrt{3}+6)?\]

myininaya (myininaya):

we are finding product right?

OpenStudy (anonymous):

\[√6(7√3 + 6)\] thats the one i still need help on

OpenStudy (anonymous):

myinaya it itwo separate equations

OpenStudy (anonymous):

distributing and simplifying

myininaya (myininaya):

\[\sqrt{6}(7 \sqrt{3}+6)\]

myininaya (myininaya):

\[\sqrt{6} \cdot 7 \sqrt{3}+\sqrt{6} \cdot 6 \]

myininaya (myininaya):

i use distributive property there

myininaya (myininaya):

a(b+c)=ab+ac

myininaya (myininaya):

\[7 \sqrt{6} \sqrt{3}+6 \sqrt{6}\] \[7 \sqrt{2 \cdot 3} \sqrt{3}+6 \sqrt{6}\]

myininaya (myininaya):

\[7 \sqrt{2} \sqrt{3} \sqrt{3}+6 \sqrt{6}\]

OpenStudy (anonymous):

is the "c" a "3"??

myininaya (myininaya):

\[7 \sqrt{2}(3)+6 \sqrt{6}=7(3) \sqrt{2}+6 \sqrt{6}=21 \sqrt{2}+6 \sqrt{6}\]

OpenStudy (anonymous):

\[ \sqrt{6}(7 \sqrt{3}+6) = 21\sqrt{2} +6\sqrt{6} \]

OpenStudy (anonymous):

Thank you guys so much!!

OpenStudy (anonymous):

Distributive property holds because they are rationals.